FIND THE MIDPOINT
p1=(5,5)
p2=(-2,6)

Answer by stanbon(75887)

FIND THE MIDPOINT
p1=(5,5)
p2=(-2,6)
-------
Let the midpoint be (x,y)
-----
x = (5+-2)/2 = 3/2
y = (5+6)/2 = 11/2
------
Cheers,
Stan H.
==========